Circular No. 7534

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 73P/SCHWASSMANN-WACHMANN 3 Recent observations suggest that three nuclear components of comet

73P are now visible: what appear to be components B and C from the observed 1995 outburst and splitting (IAUC 6246, 6274, 6301) and an apparent new component (E). Assuming that component C (T = 2001 Jan.

27) is the primary nucleus, components B and E are separated by ΔT = +0.27 and +0.74 day, respectively. Component E was observed by K. Kadota (Ageo, Japan, 0.18-m reflector + CCD) on Nov. 28.84 UT and by M. Jäger (Puchenstuben, Austria, 0.3-m reflector + Technical Pan film) on Dec. 1.19 and 2.20 — the latter indicating that it is ∼ 28′ tailward from, and ∼ 1.5–2 mag fainter than, component C. Observations by Jäger and earlier by A. Galád and P. Kolény (Modra, 0.6-m reflector + CCD) on Nov. 19.19 indicate that component B is ∼ 2.5–3 mag fainter than component C. Jäger adds that component C has a 20′ tail in p.a. 296^{o}. Total visual magnitude estimates (cf. IAUC 7523) for component C: Nov. 25.51 UT, 11.9: (A. Hale, Cloudcroft, NM, 0.2-m reflector; low altitude, zodiacal light); 28.84, 11.4 (S. Yoshida, Ibaraki, Japan, 0.25-m reflector).

SUNGRAZING COMETS Further to IAUC 7519, D. Hammer reports his measurements for sev-

eral additional Kreutz sungrazing comets found on SOHO website images obtained with the C2 coronagraph; the Nov. comets were found by M. Oates, and the Dec. comets were found by X. Leprette. The reduced observations and orbital elements by B. G. Marsden are given on the MPECs cited below. Regarding IAUC 7519 and MPEC 2000-V36, D. A. Biesecker informs us that he is not confident of the existence of comet C/1997 U6 (separated from C/1997 U5 by only 1′ in position and by 0.01 day in T).

Comet  1997 UT  α_{2000}  δ_{2000}  MPEC
C/1997 V3 Nov. 1.921  14^{h}25.^{m}7  −16^{o}09′  2000-W18
C/1997 V4  3.601  14 29.6  −16 20  2000-W18
C/1997 V5  8.564  14 53.3  −18 19  2000-W18
C/1997 V6  8.868  14 51.8  −18 08  2000-X13
C/1997 X3 Dec. 2.997  16 38.9  −23 52  2000-X13
C/1997 X4  3.080  16 39.5  −23 49  2000-X13
C/1997 X5  6.372  16 54.4  −24 14  2000-X13
C/1997 X6  10.329  17 13.9  −24 50  2000-X13
C/1997 Y1  20.889  18 02.7  −24 52  2000-X13
2000 December 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green